OpenStudy (anonymous):

If two sides of the triangle above have lengths of 5 and 6, the perimeter of the triangle could be which of the following? I. 15 II. 20 II. 22

OpenStudy (anonymous):

@pgpilot326 There can be multiple answers

OpenStudy (anonymous):

triangle above? is there a pic?

OpenStudy (anonymous):



OpenStudy (anonymous):

remember the 3rd side cannot be more than the sum of the other 2 sides. Write the equation for the perimeter and set it equal to each of the 3 values. If you solve each for the 3rd side you'll see which ones work and which don't.

OpenStudy (anonymous):

So all could work, right? @pgpilot326

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

@pgpilot326 wait, then none of them work?

OpenStudy (anonymous):

the first 2 work. the last doesn't. 5 + 6 + 4 = 15, 4 < 5 + 6 so 4 is a valid side length. 5 + 6 + 9 = 20. 9 < 5 + 6 so 9 is a valid side length. 5 + 6 + 11 = 22, 11 is not less than 5 + 6 so 11 is not a valid side length.

OpenStudy (anonymous):

OH. wow. gosh.

OpenStudy (anonymous):

make sense? in order to connect the last one would be flat.